my question is on linear equations.
Find the linear equation that models the following situation:
Your uncle Noah bought a house in 1960 for 16000 dollars. If the house value grew linearly and in 1990 the house had a value of 180000 dollars, what would the linear equation relating the value of the house(V) to the time your uncle has owned the house(t) be?

Answer by AndrewRyan(3)

The value of the house in year(y) is equal to the initial value of the house(16000) plus the product of the ValueIncreasedPerYear(x) and year(y).

(x)value increased per year = (180000 - 16000)/30
x = 164000/30
x = 16400/3

NOTE: I don't use decimals in solutions. I want the final answers to be as accurate as possible

The final equation will be
Value Of The House In Year(y) = 16000 + ValueIncreasedPerYear(x)*year(y)
value = 16000 + (16400/3)*y

checking:
In the question it is indirectly stated that the value of the house after 30 years is $180000.

value = 16000 + (16400/3)*30
value = 16000 + 164000
value = 180000